# Enhancing Robustness in Language-Driven Robotics: A Modular Approach to Failure Reduction

### v1

@@ Abstract
Recent advances in large language models (LLMs) have led to significant progress in robotics, enabling embodied agents to better understand and execute open-ended tasks. However, existing approaches using LLMs face limitations in grounding their outputs within the physical environment and aligning with the capabilities of the robot. This challenge becomes even more pronounced with smaller language models, which are more computationally efficient but less robust in task planning and execution. In this paper, we present a novel modular architecture designed to enhance the robustness of LLM-driven robotics by addressing these grounding and alignment issues. We formalize the task planning problem within a goal-conditioned POMDP framework, identify key failure modes in LLM-driven planning, and propose targeted design principles to mitigate these issues. Our architecture introduces an “expected outcomes” module to prevent mischaracterization of subgoals and a feedback mechanism to enable real-time error recovery. Experimental results, both in simulation and on physical robots, demonstrate that our approach significantly improves task success rates for pick-and-place and manipulation tasks compared to both larger LLMs and standard baselines. Through hardware experiments, we also demonstrate how our architecture can be run efficiently and locally. This work highlights the potential of smaller, locally-executable LLMs in robotics and provides a scalable, efficient solution for robust task execution. [^1] Code and experiments available at https://tinyurl.com/4h2ysybd.
@@ I Introduction
Figure 1: Overview of the proposed approach: When receiving a task from a user, a plan is formed and the expected outcomes (EOs) of each step are described. The robot then tackles the task iteratively, using skills on the environment and receiving feedback.
Recent advances in large language models (LLMs) have enabled robots to perform more complex tasks in dynamic, unstructured environments by leveraging natural language understanding. This capability has significant implications for service robotics, industrial automation, and human-robot collaboration, where robots can be directed through open-ended language instructions, making them more adaptable and user-friendly. However, despite this progress, LLM-driven robotic systems still face two critical challenges: grounding language instructions in the robot’s physical environment and aligning these instructions with the robot’s actual capabilities. These challenges are especially pronounced when using smaller, locally-executable models, which offer computational efficiency but struggle with task robustness. In this work, we propose a novel modular architecture designed to address these issues by introducing a feedback-driven alignment mechanism between language-based planning and physical execution, enabling more reliable task completion in both simulated and real-world settings.
@@
Many current methods in robotics rely on very large language models, such as GPT-3.5 (175 billion parameters) [1] or GPT-4 (more than 1 trillion parameters), to achieve reliable performance in executing complex tasks. While these models offer high accuracy, they come with significant drawbacks: their substantial computing power and energy requirements lead to high environmental and economic costs, dependence on third-party services for uptime, and potential privacy concerns due to data sharing. Despite these issues, comparatively little attention has been given to smaller language models, like LLaMA3.1 (8 billion parameters), which can be executed locally on standard hardware. These smaller models, though less powerful, still offer many of the key advantages of large language models, such as the ability to generate code and reason with natural language. However, their limited performance presents key challenges, particularly in grounding language instructions in the robot’s environment and aligning them with robot capabilities.
@@
In this paper, we address this challenge by improving the robustness and applicability of smaller language models ($<10$ billion parameters) for efficient and reliable robotic task execution. Specifically, the contributions of this paper are: (1) We formalize the problem of LLM-driven robotic task planning within a goal-conditioned POMDP framework and identify common failure modes such as subgoal mischaracterization and missing subgoals. (2) We propose design principles aimed at improving the robustness of the LLM-based planning and execution pipeline. Specifically, we introduce an “expected outcomes” module to prevent errors by inferring detailed descriptions of each plan step and a feedback mechanism to recover from execution failures. (3) Through a series of experiments, both in simulation and on physical robots, we demonstrate that our architecture significantly increases task success rates, even compared to systems using larger language models. Our pipeline is also capable of recovering from execution failures and handling tasks prompted by non-expert users.
@@ II Related works
II-A Foundation models for robotics/planning
@@
Foundation models are receiving growing attention in planning due to their compatibility with common language and open vocabularies. Some authors [2, 3] propose to use LLMs to formalize planning problems and solve them using optimal solvers, while plans inferred by LLMs can be used as a heuristic for a planner[4]. LLMs can also be used to generate behavior trees [5] or interact with scene graphs [6] when solving tasks.
@@
In the context of robotics, the common sense of LLMs can manifest as an ability to handle ambiguous instructions [7] or predict user behavior [8]. LLMs can be given access to sets of pre-defined motion primitives [7, 9], and this in turn can be used to generate datasets to train multitask policies [10]. To increase performance, authors propose to decouple the planning and code writing stages [11]. A more recent and increasingly popular trend is that of using vision-language models (VLMs) to augment the pipeline with a degree of spacial awareness. For example in COPA [12], VLMs are used to detect task-based spatial constraints and this is used for motion planning. VLMs can also be used to propose tasks grounded in the robot’s environment to collect task data [13]. Other examples of foundation models applied to robotics include generalist diffusion policies [14, 15], and related methods include using pretrained language models to encode user input for trajectory generation [16], grasping [17] or manipulation [18].
II-B Grounding and alignment in robotics
@@
When using LLMs to control robots, the disconnect between the LLM and the environment becomes an issue. To tackle this, SayCan [19] uses a value function score for each predicted action based on its feasibility and likelihood of success. SayPlan [20] and SayNav [21] for instance address the grounding problem with iterative re-planning while exploiting a 3D scene graph to align LLM output with the environment. Another approach to grounding is to leverage multi-modalities in VLM and foundation models ([22], see also above paragraph). However, most works are limited to object segmentation and not robotic manipulation. For instance, in a recent work, [23] a VLM is trained to predict affordance maps of instruction-guided manipulations.
@@
An increasingly popular technique for grounding language models is the implementation of feedback mechanisms. For example, VLM planners can detect and recover from failure [24], and human feedback, which can be coupled with uncertainty detection mechanisms [25, 26] has also been identified as a promising way of grounding LLMs [27]. For manipulation task, LLMs have also been used to infer functions measuring the success of uncertain motion primitives [10]. However, such methods rely on vision-language models or repeated human interaction.
To summarize, LLMs are gaining traction for planning and robotics applications. LLMs can be given access to a set of motion primitives, and adding vision capabilities or human feedback to the pipeline is often used to ground the model in its environment. However, most of these methods rely on very large models which can’t be run locally. This issue is even more noticeable when feedback mechanisms involve vision-language models. Human feedback is also constraining to implement, requiring a user or designer to interact with the robot. In this work, we show that most of these requirements can be relaxed, by using a smaller language model and obtaining feedback from rules verified in a simple simulator.
@@ III-A Plan execution as options in a GC-POMDP
Open-vocabulary LLM planning: We cast the problem of inferring a plan as that of finding a sequence of subgoals $\{\mathcal{G}_{i}\}_{i\in 1:n}$ such that, when all the $\mathcal{G}_{i}$’s have been reached in sequence, the task $\mathcal{G}$ is achieved. For each $i$, $\mathcal{G}_{s,i}$ is the set of states where the subgoal is fulfilled. The goal and subgoals belong to an open vocabulary. In the following, the module responsible for finding the plan is the planner.
@@
Plan execution: We consider a set of motion primitives accessible to the robot, modelled as options. As per [29], an option is a triplet $\langle\mathcal{I},\pi,\mathcal{B}\rangle$, where $\mathcal{I}$ is the initialization set, the subset of the state space from which the option’s policy is defined; $\pi$ is the policy that is followed while the option is active, which we call the intra-option policy and $\mathcal{B}$ is a termination condition. We also define $\mathcal{O}$, the set of possible outcomes of the option. This set includes both the successful outcomes $\mathcal{O}_{s}$ of the option (say for a grasping option, the object is in the robot’s gripper) and the failures $\mathcal{O}_{f}$ (the robot drops the object). We consider that the state at the end of the option is uniformly sampled from the union of the two sets. This implies that failures can randomly occur. The designer of the robotic skills doesn’t specify their exact outcomes aside from ensuring that they fulfil their purpose, meaning that more specific options can be designed, where the outcome set is a subset of a more general option’s outcome set (see Figure 2). For example, a grasping option can be parameterized to specify which part of the object is grasped.
@@
Figure 2: Options $\pi_{1}$ and $\pi_{2}$ (red) being parameterized versions of option $\pi$ (black). The outcome sets of $\pi_{1}$ and $\pi_{2}$ are subsets of the outcome set of $\pi$. For example, $\pi$ might be ”grasp mug”, with $\pi_{1}$ being ”grasp mug handle” and $\pi_{2}$ ”grasp mug body”. $O$ denotes the outcome set of $\pi$, while $O_{1}$ and $O_{2}$, respectively, denote the outcome sets of $\pi_{1}$ and $\pi_{2}$.
@@
Solving the task implies finding a sequence of options such that: (i) there are subsequences of options fulfilling each subgoal $\mathcal{G}_{i}$ sequentially; (ii) the successful outcome set of each option is in the initialization set of the following option: $\mathcal{O}_{s,i}\in\mathcal{I}_{i+1}$ (where $i$ denotes the index of the option) and (iii) option failure (which can occur randomly) is recovered from.
@@
In the following, we will say that an option is afforded by a state when this state is in the option’s initialization state. While in the context of robotics, affordances (see e.g. [30, 31]) are mostly known in the context of manipulation tasks, the concept can be used to describe the possibilities an agent has in a given environment (in our formalism, the options that can be executed from a given state).
@@
The module responsible for executing the plan steps using the options is the execution module.
III-B LLMs misalignment
@@
In this section, we assume the existence of a set of ground-truth subgoals (or GTSGs) $\{\mathcal{G}_{g,i}\}_{i\in 1:m}$ such that the transition between each pair of successive subgoals $\{\mathcal{G}_{g,i},\mathcal{G}_{g,i+1}\}$ can be achieved with one motion primitive and the plan obtained by following the GTSGs is valid. This implies that for every state fulfilling a GTSG, the following primitive is afforded by that state. We note that, since the plan is not unique, the GTSGs are more semantically complex than the plan subgoals. For example, in a pick-and-place task, the first GTSG would be states such that ”Object (a) is grasped or object (b) is grasped”.
For simplicity, in the following, we will use the notion of GTSG and of the set of states fulfilling the GTSG interchangably when there is no ambiguity.
Incorrect syntax: LLMs sometimes output code that is syntactically incorrect. This happens when primitive, location or object names are incorrect or hallucinated, or when the API’s syntax isn’t respected.
@@
Subgoal mischaracterization: Such errors arise when a subgoal isn’t properly specified by the planner. That is, the plan subgoal is not a subset of the GTSG. When executing the step, this can lead to the resulting state not fulfilling the GTSG, and not affording the remainder of the plan. In practice, such errors are difficult to recover from as the plan step appears to have been executed correctly. To avoid such errors, we propose to introduce an expected outcomes module, describing each plan step in detail. In practice, this ensures that the plan subgoal is included in the GTSG.
@@
Missing subgoal: Missing subgoals arise from discrepancies between the semantics at the plan level and at the ground-truth level. This results in the execution module’s output missing some primitives, leading to options not being afforded. Such errors can be recovered from by implementing feedback mechanisms in the system, prompting the execution module to update the code when there is an issue. In practice, we propose to implement a digital twin with simple logic rules that are used to obtain this feedback. Both types of errors are illustrated in Figure 3.
@@
Figure 3: Common errors for LLM plan execution. The GTSGs are plain blue ellipses while the plan subgoals are represented by dark blue contours. The options are in red (we omit the initialization and outcome sets of the options for legibility). Left: the plan subgoals and GTSGs are aligned. Center: mischaracterized subgoal. While the state achieves the plan subgoal, it does not fulfil the GTSG and does not afford the next primitive. Right: Missing subgoal. Missing subgoal in the plan. The plan step requires more than one option to be carried out, leading to execution errors.
@@ III-C Architecture
We introduce the main components of our architecture. For each component, we describe the key idea, inputs and outputs. .
@@
Planner: The planner receives the task, expressed in language, and outputs a series of simple steps, close to the primitive level, to tackle the task. In other words, the planner finds the set of subgoals for the agent to follow. In practice, we find that forming a plan before executing the task increases performance, facilitates recovery from errors and makes the pipeline more explainable. It also reduces the likelihood of errors, as outputting code for the whole tasks typically has a higher chance of failure. While the idea of sequentially approaching a task is related to the chain-of-thought method[32], we note that in our case the model only requires minimal documentation and examples at each stage, instead of requiring examples of the full reasoning. Exploring applications of chain-of-thought to improve the planning stage is an interesting area left for future work.
@@
Expected outcomes module: This module takes as input the task and the plan and, for each step of the plan, outputs a more detailed description of each step of the plan and its outcome. This is equivalent to providing a better characterization of the subgoal set. In practice, we also find that this module drastically reduces syntax errors in the LLM output.
@@
Execution module: This module is in charge of executing the plan, step by step, using the motion primitives available to the agent. Its main inputs are the task, plan and expected outcomes. At each time step, the module outputs a python function containing the code required to carry out the current step of the plan. In the module’s first prompt, we also provide the headers and a use example of each motion primitive, and describe the setting of the experiments (’you are in charge of a robot with one arm’). At each step, we add to the context a follow-up prompt containing the updated plan and some feedback (see next paragraph).
@@
Feedback mechanism: We implement a digital twin with simple logic rules (for example, a robot cannot grasp an object if there is already an object in the gripper). Note that this digital twin can be linked to the real world with appropriate perception module, allowing it to be updated when the world changes. If the primitives execute successfully in the digital twin, they are run in the environment. If they fail, appropriate feedback is returned to the execution module. At each time step, we also provide a reminder of the robot’s location and of the object in the gripper. The goal of this mechanism is to provide the possibility to recover from mistakes by re-attempting a failed step, whether that failure is due to erroneous syntax, primitive failure or planning issues. In our experiments, we focus on simple feedback mechanisms that can be implemented without VLMs to enforce transferability to the real world. In Sections IV and V, we show that a few (two or three per skill) simple checks drastically robustify task execution.
Figure 4: Planning and execution architecture. The task, plan, expected outcomes and system feedback are formulated in language, while the instructions that are passed to the system are in Python.
@@
Implementation details: We implement our architecture using ROS. The planner and expected outcomes module are in the same program, with both prompting separate instances of LLaMA3.1 in turn when the task is published. When ready, the plan and expected outcomes are published by this program. The second program contains the execution module. When both the plan and expected outcomes have been received, the execution module receives its initial prompt. This prompt contains the task, plan, skills, a description of the setting (robot, number of arms) and environment (locations, objects). When appropriate (in our ablation study, we benchmark our full architecture with variants where the EO module is inactive, see Section IV-C), the expected outcomes are also added to this prompt. When the execution module outputs some code, the code is published and executed in the digital twin, and the execution module receives feedback: either ’Done’ if the code runs correctly or simple feedback when errors occur. The feedback also includes the robot’s position and the object in the gripper. In the experiments where the feedback pipeline isn’t active, the feedback is always ’Done’. If the step was successful, the current step is removed from the plan. Finally, the LLM is reprompted (by adding to the context), with the feedback, the plan and, if applicable, the current expected outcome, closing the loop.

### v2

@@ Abstract
Recent advances in large language models (LLMs) have led to significant progress in robotics, enabling embodied agents to understand and execute open-ended tasks. However, existing LLM-based approaches face limitations in grounding their outputs within the physical environment and aligning with the capabilities of the robot. While fine-tuning is an attractive approach to addressing these issues, the required data can be expensive to collect, especially when using very large language models. Smaller language models, while more computationally efficient, are less robust in task planning and execution, leading to a difficult trade-off between performance and tractability. In this paper, we present a novel, modular architecture designed to enhance the robustness of locally-executable LLMs in the context of robotics by addressing these grounding and alignment issues. We formalize the task planning problem within a goal-conditioned POMDP framework, identify key failure modes in LLM-driven planning, and propose targeted design principles to mitigate these issues. Our architecture introduces an “expected outcomes” module to prevent mischaracterization of subgoals and a feedback mechanism to enable real-time error recovery. Experimental results, both in simulation and on physical robots, demonstrate that our approach leads to significant improvements in success rates for pick-and-place and manipulation tasks, surpassing baselines using larger models. Through hardware experiments, we also demonstrate how our architecture can be run efficiently and locally. This work highlights the potential of smaller, locally-executable LLMs in robotics and provides a scalable, efficient solution for robust task execution and data collection. [^1] Code available at https://tinyurl.com/4h2ysybd.
{strip}[Uncaptioned image]
Figure 1: Overview of the proposed approach: When receiving instructions from a user, a plan and the expected outcomes (EOs) of each step are inferred. The robot then tackles the task iteratively, using motion primitives and receiving feedback from a simulation, before execution in the environment. See Section III for notations and formalism.
@@ I Introduction
Due to their natural language understanding and code writing abilities, large language models (LLMs) are an attractive tool for robots evolving in open environments. This is especially true in service robotics, industrial automation, and human-robot collaboration, where robots can be directed through open-ended language instructions, making them more adaptable and user-friendly. However, LLM-driven robotic systems still face two critical challenges: grounding language instructions in the robot’s physical environment and aligning these instructions with the robot’s actual capabilities. Many current methods in robotics rely on very large language models, such as GPT-3.5 ($175$ billion parameters) [1] or GPT-4o (more than $1.7$ trillion parameters), to achieve reliable performance when executing complex tasks. While these models offer high accuracy, they come with significant drawbacks: their substantial computing power and energy requirements lead to high environmental and economic costs, dependence on third-party services for uptime, and potential privacy concerns due to data sharing. Despite these issues, comparatively little attention has been given to smaller language models, like LLaMA3.1 (8 billion parameters), which can be executed locally on standard hardware. These smaller models, though less powerful, still offer many of the key advantages of large language models, such as the ability to generate code and reason with natural language. However, their limited performance exacerbates the grounding and alignment problems. While fine-tuning is often proposed to address such challenges, the design of scalable, zero-shot-compatible data collection architectures remains a challenge.
@@
In this paper, we address this challenge by improving the robustness and applicability of smaller language models ($<10$ billion parameters) for efficient and reliable robotic task execution. Specifically, our contributions are: (1) We formalize the problem of LLM-driven robotic task planning within a goal-conditioned POMDP framework and identify two failure modes: subgoal mischaracterization and missing subgoals. (2) We propose design principles aimed at improving the robustness of the LLM-based planning and execution pipeline. Specifically, we introduce an “expected outcomes” module to prevent errors by inferring detailed descriptions of each plan step and a feedback mechanism to recover from execution failures. (3) Through a series of experiments, both in simulation and on physical robots, we demonstrate that our architecture significantly increases task success rates, even outperforming systems using larger language models while incurring significantly lower computational costs. Our pipeline is also capable of handling instructions given by non-expert users, and, by splitting task execution in several plan steps, is more explainable and can recover from execution failures. We also (4) test one of the dense, distilled versions of Deepseek (Deepseek-r1:7b), exploring its applicability to robotics use cases.
@@ II Related works
II-A Foundation models for robotics and planning
@@
Foundation models are receiving growing attention in planning due to their compatibility with natural language and open vocabularies. Some authors [2, 3] propose to use LLMs to formalize planning problems and solve them using optimal solvers, and plans inferred by LLMs can be used as a heuristic for a planner[4]. LLMs can also be used to generate behavior trees [5] or interact with scene graphs [6] when solving tasks.
@@
In the context of robotics, the common sense of LLMs can manifest as an ability to handle ambiguous instructions [7] or predict user behavior [8]. LLMs can be given access to sets of pre-defined motion primitives [7, 9], and this in turn can be used to generate datasets to train multitask policies [10]. To increase performance, authors propose to decouple the planning and code writing stages [11]. LLM-inferred plans can also be used as a basis for reward function writing [12]. An increasingly popular trend is that of using vision-language models (VLMs) to augment the pipeline with a degree of spatial awareness. For example in COPA [13], VLMs are used to detect task-based spatial constraints and this is used for motion planning. VLMs can also be used to propose tasks grounded in the robot’s environment to collect task data [14].
II-B VLA models
Vision-language-action (VLA) models are a rapidly emerging paradigm in robotics. The key idea of VLAs is to directly predict actions from the task statement and robot camera(s) output[15]. While early VLA models were trained from scratch, using language models as a backbone has been gaining traction, with the goal of leveraging such models’ general knowledge [16], or even retaining multimodal understanding abilities [17]. VLAs have shown impressive results on challenging manipulation tasks, but generalization remains a challenge and they often rely on fine-tuning[18], requiring large datasets.
II-C Grounding and alignment in robotics
@@
In robotics, the disconnect between the LLM and the environment is an issue. To tackle this, SayCan [19] uses a value function score for each predicted action based on its feasibility and likelihood of success. SayPlan [20] and SayNav [21] address the grounding problem with iterative re-planning, exploiting a 3D scene graph. Another approach to grounding is to leverage multi-modalities in VLMs and VLAs ([22], see also above).
@@
An increasingly popular technique for grounding language models is the implementation of feedback mechanisms. For example, VLM planners can detect and recover from failure [23], and human feedback, which can be coupled with uncertainty detection mechanisms [24, 25] has also been identified as a promising way of grounding LLMs [26]. For manipulation task, LLMs have also been used to infer functions measuring the success of uncertain motion primitives [10]. Training models [27, 28] with domain-specific data is also gaining traction, but collecting enough high-quality data remains a challenge.
Studies similar to ours include MALMM [29], where a multi-agent architecture plans and executes a task. However, the architecture doesn’t consider explicit expected outcome characterization. Further, it relies on the LLM agents’ ability to properly parametrize geometric motion primitives, which heavily relies on the high performance of GPT-4o, leading to scalability issues. Another work [30] envisions a two-level error management pipeline. However, at the low level, error correction mostly relies on automated fallbacks, while we propose using the execution LLM to handle errors, and using a preprocessing step to prevent them. We finally note that, by design, our pipeline is compatible with high-level plan adaptation techniques.
Overall in robotics, foundation models still suffer from issues with alignment and grounding, and most current methods rely on the impressive capabilities of very large models. In this work, we show that feedback mechanisms and explicit subgoal characterization can yield significant increases in performance when using smaller LLMs, and lead to a more explainable, scalable and less computationally demanding system.
@@ III-A Plan execution as options in a GC-POMDP
Open-vocabulary LLM planning: Planning is the problem of finding a sequence of plan steps $\{P_{i}\}_{i\in 1:n}$, or, equivalently, subgoals $\{\mathcal{G}_{i}\}_{i\in 1:n}$ such that, when all the $\mathcal{G}_{i}$’s have been reached in sequence, the task $\mathcal{G}$ is achieved.
@@
Plan execution: We consider a set of motion primitives accessible to the robot, modelled as options. As per [32], an option is a triplet $\langle\mathcal{I},\pi,\mathcal{B}\rangle$, where $\mathcal{I}$ is the initialization set, the subset of the state space from which the option’s policy is defined; $\pi$ is the policy that is followed while the option is active, which we call the intra-option policy and $\mathcal{B}$ is a termination condition. We also define $\mathcal{O}$, the set of possible outcomes of the option. This set includes both the successful outcomes $\mathcal{O}_{s}$ of the option (say for a grasping option, the object is in the robot’s gripper) and the failures $\mathcal{O}_{f}$ (the robot drops the object). We consider that the state at the end of the option is uniformly sampled from the union of the two sets. This implies that failures can randomly occur. The designer of the robotic skills doesn’t specify their exact outcomes aside from ensuring that they fulfil their purpose, meaning that parametrized options can be designed, where the outcome set is a subset of a more general option’s outcome set (see Figure 2). For example, a grasping option can be parameterized to specify which part of the object is grasped.
@@
Figure 2: Options $\pi_{1}$ and $\pi_{2}$ (red) being parameterized versions of option $\pi$ (blue). The outcome sets of $\pi_{1}$ and $\pi_{2}$ are subsets of the outcome set of $\pi$. For example, $\pi$ might be ”grasp mug”, with $\pi_{1}$ being ”grasp mug handle” and $\pi_{2}$ ”grasp mug body”. $O$ denotes the outcome set of $\pi$, while $O_{1}$ and $O_{2}$, respectively, denote the outcome sets of $\pi_{1}$ and $\pi_{2}$.
@@
Solving the task implies finding a sequence of options such that: (i) the options fulfil each subgoal $\mathcal{G}_{i}$ sequentially; (ii) the successful outcome set of each option is in the initialization set of the following option: $\mathcal{O}_{s,i}\subset\mathcal{I}_{i+1}$ (where $i$ denotes the index of the option) and (iii) option failure (which can occur randomly) is recovered from.
@@
In the following, we will say that an option is afforded by a state when this state is in the option’s initialization state. While in robotics, affordances (see e.g. [33, 34]) are mostly known in the context of manipulation tasks, the concept can be used to describe the possibilities an agent has in a given environment (in our formalism, the options that can be executed from a state).
@@
The module responsible for executing the plan steps using the options is the execution module. In this work, the execution module outputs code snippets $C$ in which the options are called through python functions.
III-B LLM planning failures
@@
In this section, we assume the existence of a set of ground-truth subgoals (or GTSGs) $\{\mathcal{G}_{g,i}\}_{i\in 1:m}$ such that the transition between each pair of successive subgoals $\{\mathcal{G}_{g,i},\mathcal{G}_{g,i+1}\}$ can be achieved with one motion primitive and the plan obtained by following the GTSGs is valid. After surveying LLM planning, we identifed two common failure modes (see also Figure 3):
@@
Subgoal mischaracterization: Such errors arise when a subgoal isn’t properly specified by the planner. That is, the plan subgoal is not a subset of the GTSG. When executing the step, this can lead to the resulting state not fulfilling the GTSG, and not affording the remainder of the plan. In practice, such errors arise from a lack of precision or missing information in the plan step, and are difficult to recover from as the plan step appears to have been executed correctly. To avoid such errors, we propose to introduce an expected outcomes module, explicitly characterizing the plan subgoal and ensuring that it is a subset of the GTSG.
@@
Missing subgoal: Missing subgoals arise from discrepancies between the semantics at the plan level and at the ground-truth level. This leads to plans that are correct, but cannot be achieved by simply mapping primitives to each plan step. This results in the execution module’s output missing some primitives (for example not moving before trying to grasp an object), leading to options not being afforded. Such errors can be recovered from by implementing feedback mechanisms in the system, prompting the execution module to update the code by adding the missing primitives, compensating for the missing plan step. See Section III-C for more details on these mechanisms.
@@
Figure 3: Common errors for LLM plan execution. The GTSGs are plain blue ellipses while the plan subgoals are represented by dark blue contours. The options are in red (we omit the initialization and outcome sets of the options for legibility). Left: the plan subgoals and GTSGs are aligned. Center: Mischaracterized subgoal: while the state achieves the plan subgoal, it does not fulfil the GTSG and does not afford the next primitive. Right: Missing subgoal in the plan: the plan step requires more than one option to be carried out, leading to execution errors.
@@ III-C Architecture
We introduce the main components of our architecture. For each component, we describe the key idea, inputs and outputs. An illustration of the complete architecture can be found in Figure 1 and prompts for each module can be found in the appendix.
@@
Planner: The planning module receives the task instruction, expressed in language, and outputs a series of simple steps, close to the primitive level, achieving the task. In other words, the planner finds a set of subgoals for the agent to follow. In practice, we find that forming a plan before executing the task increases performance, facilitates recovery from errors and makes LLM-based task execution more explainable. It also reduces the likelihood of errors, as outputting code for the whole tasks typically has a higher chance of failure. While the idea of sequentially approaching a task is inspired by the chain-of-thought method[35], we note that in our case the model responsible for planning is used in a zero-shot manner, instead of requiring examples of full reasoning chains.
@@
Expected outcomes module: This module takes as input the instruction and the plan and outputs the expected outcome of each step of the plan, in terms of robot and world state. This is to provide a more precise characterization of each subgoal, resulting in a more complete plan. We also find that this module also improves the code quality of downstream modules, reducing hallucinations in code output.
@@
Execution module: This module is in charge of executing the plan, step by step, using the motion primitives available to the agent. Its main inputs are the instruction, plan and expected outcomes. At each time step, the module outputs a python code snippet $C$ containing the code required to carry out the current step of the plan. In the module’s first prompt, we also provide the headers and a use example of each motion primitive, and describe the setting of the experiments (’you are in charge of a robot with one arm’). At each step, we add to the context a follow-up prompt containing the updated plan and some feedback (see next paragraph).
@@
Feedback mechanism: We implement a symbolic simulation with simple logic rules (for example, a robot cannot grasp an object if there is already an object in the gripper or if it is not at the same location as the object). Note that this simulation can be linked to the real world using perception modules, allowing for the feedback mechanism to be grounded. If the primitives execute successfully in the digital twin, they are executed in the environment. If they fail, appropriate feedback is returned to the execution module. At each time step, we also provide a reminder of the robot’s location and of the object in the gripper. The goal of this mechanism is to provide the possibility to recover from code writing mistakes, primitive failure or planning issues by re-attempting a failed step. In Sections IV and V, we show that a few (two or three per skill) simple checks drastically robustify task execution.
@@
Implementation details: We implement our architecture using ROS. The planner and expected outcomes module are in the same program, with both prompting separate instances of the LLM in turn when the task is published. When ready, the plan and expected outcomes are published by this program. The second program contains the execution module. When both the plan and expected outcomes have been received, the execution module receives its initial prompt. This prompt contains the task, plan, skills, a description of the setting (robot, number of arms) and environment (locations, objects). When appropriate (in our ablation study, we benchmark our full architecture with variants where the EO module is inactive, see Section IV-C), the expected outcomes are also added to this prompt. When the execution module outputs some code, the code is published and executed in the digital twin, and the execution module receives feedback: either ’Done’ if the code runs correctly or simple feedback when errors occur. The feedback also includes the robot’s position and the object in the gripper. In the experiments where the feedback pipeline isn’t active, the feedback is always ’Done’. If the step was successful, the current step is removed from the plan and the code is executed in the environment. Finally, the LLM is reprompted (by adding to the context), with the feedback, the plan and, if applicable, the current expected outcome, closing the loop.